OpenStudy (anonymous):

Solve 5x2 = -30x - 65.

OpenStudy (anonymous):

Help!!!

OpenStudy (anonymous):

This isn't too hard, just isolate x to one side.

OpenStudy (anonymous):

\[ 5x^2 = -30x - 65\] start with \[5x^2+30x+65=0\] then maybe factor

OpenStudy (anonymous):

Satellite explained it. And wo, I've never seen a 100 as a SmartScore.

OpenStudy (anonymous):

even easier if you divide by \(5\) and start with \[x^2+6x+13=0\] which i now see does not factor, hmm guess you need the quadratic formula

OpenStudy (anonymous):

and in fact there are no real solutions, only two complex ones

OpenStudy (anonymous):

x = -3 ± 2i x = -3 ± 4i x = -6 ± 2i x = -6 ± 4i

OpenStudy (anonymous):

yes, it is one of those definitely complex

OpenStudy (anonymous):

is it a or c\

OpenStudy (anonymous):

easiest to solve in this case by completing the square, because the middle term is \(6x\) you have \[x^2+6x+13=0\\ x^2+6x=-13\\ (x+3)^2=-13+3^2\\ (x+3)^2=-4\\ x+3=\pm\sqrt{-4}\\ x+3=\pm2i\\x=-3\pm2i\]

OpenStudy (anonymous):

Solve -2x2 +3x - 9 = 0.